### v1

@@ A.1 Proofs
In order to prove that $\sum_{i\in\tilde{I}}\bar{w}_{\tilde{R}_{i}}\tilde{x}_{\tilde{R}_{i}}^{T}\hat{x}_{A}>\sum_{i\in I}\bar{w}_{R_{i}}x_{R_{i}}^{T}\hat{x}_{A}$ it is sufficient to follow the same steps of used in proposition 3.2 to show that $\bar{x}^{\prime}-\underaccent{\bar}{x}^{\prime}>\bar{x}-\underaccent{\bar}{x}$. In particular, the proof follows from the one of proposition 3.2 if we impose (i) $I^{\prime}=\tilde{I}$; (ii) for each $i\in\tilde{I}$, $\bar{w}^{\prime}_{i}=\underaccent{\bar}{w}_{N-i}^{\prime}=\bar{w}_{\tilde{R}_{i}}$, $x^{\prime}_{i}=-x^{\prime}_{N-i}=\tilde{x}_{\tilde{R}_{i}}^{T}\hat{x}_{A}$, $s^{\prime}_{i}=s^{\prime}_{N-i}=\tilde{s}_{\tilde{R}_{i}}$; and (iii) for each $i\in I$, $\bar{w}_{i}=\underaccent{\bar}{w}_{N-i}=\bar{w}_{R_{i}}$, $x_{i}=-x_{N-i}=x_{R_{i}}^{T}\hat{x}_{A}$, and $s_{i}=s_{N-i}=s_{R_{i}}$.
As for the second part of the proof, we start by noting that the inductive step of the proof of lemma A.2 still holds at when we consider distribution $(\tilde{\mathbf{x}},\tilde{\mathbf{s}})$ and initiate the best response dynamics at $\hat{x}$: by using the exact same steps, one can indeed show that if the inequality 7 holds at some odd step $n$ of the best response sequence, then it must hold at step $n+2$, and 8 must hold at $n+1$ of the best response sequence. Hence, it is sufficient to show that $||x_{A1}||^{2}-||x_{B1}||^{2}>\frac{1}{2}||\hat{x}_{A}-\hat{x}_{B}||^{2}+2x_{A1}^{T}x_{B1}$, that is, that inequality 7 indeed holds at $n=1$.

### v2

@@ A.1 Proofs
In order to prove that $\sum_{i\in\tilde{I}}\bar{w}_{\tilde{R}_{i}}\tilde{x}_{\tilde{R}_{i}}^{T}\hat{x}_{A}>\sum_{i\in I}\bar{w}_{R_{i}}x_{R_{i}}^{T}\hat{x}_{A}$ it is sufficient to follow the same steps of used in proposition 3.2 to show that $\bar{x}^{\prime}-\underaccent{\bar}{x}^{\prime}>\bar{x}-\underaccent{\bar}{x}$. In particular, the proof follows from the one of proposition 3.2 if we impose (i) $I^{\prime}=\tilde{I}$; (ii) for each $i\in\tilde{I}$, $\bar{w}^{\prime}_{i}=\underaccent{\bar}{w}_{N+1-i}^{\prime}=\bar{w}_{\tilde{R}_{i}}$, $x^{\prime}_{i}=-x^{\prime}_{N+1-i}=\tilde{x}_{\tilde{R}_{i}}^{T}\hat{x}_{A}$, $s^{\prime}_{i}=s^{\prime}_{N+1-i}=\tilde{s}_{\tilde{R}_{i}}$; and (iii) for each $i\in I$, $\bar{w}_{i}=\underaccent{\bar}{w}_{N+1-i}=\bar{w}_{R_{i}}$, $x_{i}=-x_{N+1-i}=x_{R_{i}}^{T}\hat{x}_{A}$, and $s_{i}=s_{N+1-i}=s_{R_{i}}$.
As for the second part of the proof, we start by noting that the inductive step of the proof of lemma A.2 still holds when we consider distribution $(\tilde{\mathbf{x}},\tilde{\mathbf{s}})$ and initiate the best response dynamics at $\hat{x}$: by using the exact same steps, one can indeed show that if the inequality 7 holds at some odd step $n$ of the best response sequence, then it must hold at step $n+2$, and 8 must hold at $n+1$ of the best response sequence. Hence, it is sufficient to show that $||x_{A1}||^{2}-||x_{B1}||^{2}>\frac{1}{2}||\hat{x}_{A}-\hat{x}_{B}||^{2}+2x_{A1}^{T}x_{B1}$, that is, that inequality 7 indeed holds at $n=1$.